OpenStudy (anonymous):

A rancher wants to fence in an area of 1000000 square feet in a rectangular field and then divide it in half with a fence down the middle parallel to one side. What is the shortest length of fence that the rancher can use?

OpenStudy (anonymous):

Okay. Let L represent the length of the field. Let W represent the width of the field. Let F represent the length of fence needed. L*W=2400000 2L+3W=F Both length sides+both width sides+ fence across the width. Let's look at the first equation again: L*W=2400000 W=(2400000/L) Now we can substitute (2400000/L) for W in the second equation. 2L+3(2400000/L)=F 2L+7200000/L=F Now we have a function. We can input any length, and we'll know length needed. But how can we find the length needed so that F is as small as it can be? If you know derivatives, there's an exact way. The next easiest way to do it is to graph. Manner 1: Derivative Set the derivative to zero...when the length stops going down and starts coming back up, it'll be at it's lowest point. 2L+7200000L^-1=F F'=2-7200000L^-2 0=2-7200000L^-2 7200000/(L^2)=2 7200000=2L^2 3600000=L^2 L=1897.366 Now, we know that L=2400000/W so 1897.366=2400000/W W=2400000/1897.366 W=1264.91106 feet. Now, 2L+3W=2(1897.366)+3(1264.91106) Minimum fencing needed: 7589.465 Method 2: Graphing. Graph y=2x+7200000/x (y=fence size, x=length of side) to to the "bottom" or the graph. Zoom in really far. As we zoom in closer and closer, we find that x=1897.366 and y=7589.465. Thus, the minimum amount of fencing needed is 7589.465 feet.

OpenStudy (anonymous):

Hope this helped! Have a great day! @kdancer03

OpenStudy (anonymous):

where did the 240000

OpenStudy (anonymous):

where did the 2400000 come from?

OpenStudy (anonymous):

help would be much appreciated